Grade 9Mechanics

if gravel if falling on a conveyor belt at the rate of 0.5kg/sec.............how much extra force ( in N) is needed so as to keep the belt moving at 2m/s.................pls EXPLAIN EACH STEP!!!

Shivam Bhagat
14 Years agoGrade 9

1 Answer

Chetan Mandayam Nayakar
14 Years ago

F=d/dt(mv)=mdv/dt+vdm/dt=0.5*2=1 N